OpenStudy (anonymous):

Factor the polynomial by grouping. 5x^4-2x^2+20x^2-8

OpenStudy (anonymous):

5x^4+18x^2-8

OpenStudy (anonymous):

=5x^4+(-2+20)x^2-8 =5x^4+(18)x^2-8 =5x^4+18x^2-8 =(x^2+4)(5x^2-2) (gave it a shot!!) :)

OpenStudy (anonymous):

Thank you! :)